OpenStudy (anonymous):

Find the general solution of the following second order differential equation

OpenStudy (anonymous):

\[\frac{ d ^{2}y }{ dx ^{2} } + \frac{ dy }{ dx } = x ^{2}\]

OpenStudy (perl):

you need to first solve the homogenous equation

OpenStudy (perl):

The general solution is: y(x) = y_c(x)+y_p(x) = x^3/3-x^2+2 x+c_1 e^(-x)+c_2

OpenStudy (anonymous):

how do you solve the homogenous equation? the x^2 is what's putting me off, if it was x i would easily be able to solve it

OpenStudy (anonymous):

@perl

OpenStudy (perl):

to solve the homogenous, you have y ' ' + y ' = 0

OpenStudy (perl):

the characteristic equation is r^2 + r = 0 r ( r + 1) = 0 r = 0, r = -1 y = c_1 * e^ (0*x) + c2 * e^(-1*x)

OpenStudy (perl):

that simplifies to y = c_1 * 1 + c2 * e^ -x

OpenStudy (anonymous):

thanks so much @perl it all makes sense now! just having one of those days haha